# Synergistic eigenanalysis of covariance and Hessian matrices for enhanced binary classification on health datasets

### v2

@@ Abstract
Covariance and Hessian matrices have been analyzed separately in the literature for classification problems. However, integrating these matrices has the potential to enhance their combined power in improving classification performance. We present a novel approach that combines the eigenanalysis of a covariance matrix evaluated on a training set with a Hessian matrix evaluated on a deep learning model to achieve optimal class separability in binary classification tasks. Our approach is substantiated by formal proofs that establish its capability to maximize between-class mean distance and minimize within-class variances, particularly under ideal data conditions such as isotropy around class means and dominant leading eigenvalues. By projecting data into the combined space of the most relevant eigendirections from both matrices, we achieve optimal class separability as per the linear discriminant analysis (LDA) criteria. Empirical validation across neural and health datasets consistently supports our theoretical framework and demonstrates that our method outperforms established methods. Our method stands out by addressing both LDA criteria, unlike PCA and the Hessian method, which predominantly emphasize one criterion each. This comprehensive approach captures intricate patterns and relationships, enhancing classification performance. Furthermore, through the utilization of both LDA criteria, our method outperforms LDA itself by leveraging higher-dimensional feature spaces, in accordance with Cover’s theorem, which favors linear separability in higher dimensions. Our method also surpasses kernel-based methods and manifold learning techniques in performance. Additionally, our approach sheds light on complex DNN decision-making, rendering them comprehensible within a 2D space.
@@ 2.1 Approach: combining eigenanalysis of covariance and Hessian matrices
    Covariance matrix eigenanalysis: A covariance matrix $\text{Cov}(\boldsymbol{\theta})$ is a $D\times D$ matrix, where $D$ represents the dimensionality of the predictor attributes. Each element of the covariance matrix reflects the covariance between the two predictor attributes $\theta_{1}$ and $\theta_{2}$:
    $\text{{Cov}}(\theta_{1},\theta_{2})=\frac{1}{n-1}\sum_{i=1}^{n}(\theta_{1i}-\bar{\theta}_{1})(\theta_{2i}-\bar{\theta}_{2})$ (1)
    where $\theta_{1i}$ and $\theta_{2i}$ are the corresponding observations of these predictor attributes for the $i$-th data instance, $\bar{\theta}_{1}$ and $\bar{\theta}_{2}$ are the sample means of the predictor attributes $\theta_{1}$ and $\theta_{2}$, respectively, and $n$ is the number of data instances or observations.
    Performing eigenanalysis on the covariance matrix $\text{Cov}(\boldsymbol{\theta})$, we obtain eigenvalues $\lambda_{i}$ and corresponding eigenvectors $\mathbf{v}_{i}$. The leading eigenvector $\mathbf{v}_{1}$ associated with the largest eigenvalue $\lambda_{1}$, as captured in the eigen-equation $\text{Cov}(\boldsymbol{\theta})\cdot\mathbf{v}_{1}=\lambda_{1}\cdot\mathbf{v}_{1}$, represents the principal direction with the highest variance.
@@
2.2.1 Theorem 1: Maximizing covariance for maximizing squared between-class mean distance.
Consider two sets of 1D data points representing two classes, denoted as $C_{1}$ and $C_{2}$, each consisting of $n$ samples. The data in $C_{1}$ and $C_{2}$ follow the same underlying distribution centered around their respective means, $m_{1}$ and $m_{2}$. Here, $C_{2}$ can be understood as a reflection of $C_{1}$ with the axis of reflection positioned at the overall mean, denoted as $m$. Furthermore, the variances of $C_{1}$ and $C_{2}$ are equal, denoted as $\sigma_{1}^{2}=\sigma_{2}^{2}=\sigma_{w}^{2}$, and the combined data from $C_{1}$ and $C_{2}$ has a variance of $\sigma^{2}$. The between-class mean distance, denoted as $d$, represents the separation between the means of $C_{1}$ and $C_{2}$. We establish the following relationship:
$\sigma^{2}=\frac{d^{2}}{4(1-\lambda)}$ (6)
where $\lambda=\frac{\sigma_{w}^{2}}{\sigma^{2}}$ signifies a constant between 0 and 1 that reflects the distribution of the original data when $C_{1}$ and $C_{2}$ are considered as projected representations.
Proof sketch:
@@
    Simplify the expression for the combined data variance $\sigma^{2}$ to $\sigma_{w}^{2}+\frac{1}{4}d^{2}$.
@@
    Apply the Variance Ratio Preservation Theorem for Projection onto a Vector to relate $\sigma_{w}^{2}$ and $\sigma^{2}$ through $\lambda$.
@@
In the context of multidimensional data, extending Theorem 1 involves considering data points as $D$-dimensional vectors, where the dataset consists of two subsets, $C_{1}=\{\vec{x}_{i}\in\mathbf{R}^{D}\mid i=1,2,\ldots,n\}$ with covariance $Cov_{1}$, and $C_{2}=\{\vec{y}_{j}\in\mathbf{R}^{D}\mid j=1,2,\ldots,n\}$ with covariance $Cov_{2}$. The theorem, however, is supposed to hold for the projection of the data onto an arbitrary 1D subspace of the full D-dimensional data space, and to imply that maximizing $\sigma^{2}$ by choosing the direction of this subspace appropriately will also be the direction that maximizes the difference between the projected sample means. A general case in which it holds is when the individual distributions are isotropic about their means, meaning the covariances $Cov_{1}$ and $Cov_{2}$ are proportional to the identity matrix in $\mathrm{D}$ dimensions as mathematically shown in B.
2.2.2 Theorem 2: Maximizing Hessian for minimizing within-class variance
@@
$\mathrm{H}_{\theta}=\frac{1}{\sigma_{\text{post}}^{2}}.$ (7)
Proof sketch:
@@ 3 Ideal dataset conditions for the optimal performance of the proposed method
      Normalized data: Normalizing the data, typically through techniques such as z-score normalization, ensures that each attribute has a mean of zero and a standard deviation of one. This process equalizes variance in all directions, which approximates the isotropic condition.
@@
      Low correlation between attributes: Datasets with low correlation between attributes produce covariance matrices with minimized off-diagonal elements. This condition approximates isotropy, which is characterized by covariance matrices being proportional to the identity matrix.
@@
    The presence of a dominat first eigenvalue can be assessed from the eigenspectra of the two matrices, which display the eigenvalues on a log scale. A dominant first eigenvalue is characterized by the eigenvalues being spread over many orders of magnitude, often with approximately uniform spacing between them. When a dominant first eigenvalue is present in the covariance matrix, it suggests a dominant underlying structure or trend in the data. When this occurs in the Hessian matrix, it characterizes a ”sloppy” model [53, 54, 55, 56, 57, 58, 59], which has a few stiff directions (significant eigenvalues) and many sloppy directions (small eigenvalues).
@@ 4 Complexity analysis
In assessing the computational demands of our proposed method, we employ big O notation to describe the worst-case time complexity. The overall time complexity of our method encompasses several tasks, including covariance and Hessian matrix computations ($\mathcal{O}(N\cdot F^{2})$), eigenanalysis of these matrices ($\mathcal{O}(F^{3})$), selection of eigenvectors corresponding to leading eigenvalues ($\mathcal{O}(1)$), and data projection into the combined space ($\mathcal{O}(N\cdot F)$). Here, $N$ signifies the number of data points in the training set, while $F$ represents the number of features per data point. It is essential to note that this analysis considers our method’s computational demands under the assumption that a pre-existing DNN is in place.
@@ 5.1 Assessed dimensionality reduction techniques
In our evaluation, we compare nine distinct dimensionality reduction and data projection techniques, each offering unique insights: principal component analysis (PCA), kernel PCA (KPCA) [60], Hessian, uniform manifold approximation and projection (UMAP) [61], locally linear embedding (LLE) [62], linear optimal low-rank projection (LOL) [63], linear discriminant analysis (LDA), kernel discriminant analysis (KDA) [64], and the proposed approach. PCA involves projection onto the primary two covariance eigenvectors, i.e., it applies KPCA with a linear kernel. Similarly, LDA employs KDA with a linear kernel. The Hessian method projects data onto the leading two Hessian eigenvectors. Notably, for both KPCA and KDA, the best kernels (linear, polynomial, RBF, sigmoid, or cosine similarity) and their associated parameters (kernel coefficient or degree) are determined using grid search tailored to each dataset. The proposed method combines the most relevant eigenvectors derived from both the covariance and Hessian matrices, as specified by Eqs(4, 5).
@@ 6 Illustrative case study: WBCD dataset
Figure 3: Pairplot of the first four predictor attributes in the WBCD dataset, colored by class labels. This pairplot presents only four of the 30 predictor attributes, projecting the 30-dimensional data into 1D and 2D spaces between individual attributes and attribute pairs. The visualizations show that these projections do not provide good separability between the two classes.
Our proposed method leverages the strengths of both the covariance matrix and the Hessian matrix to achieve optimal class separability. Figure 4 presents the heatmaps of the covariance and Hessian matrices computed from the WBCD dataset. The covariance matrix, computed unsupervisedly from the training set using Eq. 1, is a square and symmetric matrix that captures the statistical spread and relationships between predictor attributes. The Hessian matrix, evaluated on a deep learning model trained on the same training set using Eq. 3, is another square and symmetric matrix that reflects the sensitivity and curvature of the loss function with respect to the model parameters. The DNN test scores are as follows: accuracy: 0.9883, F1 score: 0.9841, AUC ROC: 0.9917, Cohen Kappa score: 0.9749, geometric mean score: 0.9874.
Figure 4: Heatmaps of the covariance matrix and the Hessian matrix for the WBCD dataset. The covariance matrix is computed unsupervisedly from the WBCD training set using Eq. 1. The Hessian matrix is evaluated on a deep learning model trained on the same training set using Eq. 3.
By performing eigenanalysis on the covariance and Hessian matrices, we obtain leading eigenvectors that represent the principal directions of variance and the sharpest curvature, respectively. These eigendirections are crucial for projecting the high-dimensional data into a 2D space that enhances class separability by maximizing the between-class mean distance and minimizing the within-class variances, as described in Theorem 1 and Theorem 2. We project the data into the combined space of the two eigendirections using Eq. 4 and Eq. 5. Figure 5 visually demonstrates the effectiveness of the proposed method by showing the 2D projections of the WBCD dataset. The training set projection highlights the observed visual separation achieved on the training data, while the test set projection confirms good visual separability for new unseen data, underscoring the method’s potential to generalize well. The visual separability observed here will be quantified later using SVM (Figures 7 and 8).
@@
While Figures 5 illustrate the visual separations achieved by the proposed method, Figures 7 and 8 quantify their degrees of separation. These figures present the classification performance using various dimensionality reduction methods, including PCA, the Hessian method, LDA, UMAP, and the proposed method, evaluated with linear SVM on the WBCD dataset. Each subplot shows the separating SVM hyperplane along with the corresponding F-score, providing a quantitative measure of the classification performance. Figure 7 highlights the separation achieved on the training data, where UMAP achieves the highest performance with perfect separation, while the proposed method demonstrates strong performance, surpassing other methods. Figure 8 shows the performance on new unseen test data, where the proposed method outperforms all other methods, including UMAP, demonstrating its superior generalization capability.

### v3

@@ Abstract
Covariance and Hessian matrices have been analyzed separately in the literature for classification problems. However, integrating these matrices has the potential to enhance their combined power in improving classification performance. We present a novel approach that combines the eigenanalysis of a covariance matrix evaluated on a training set with a Hessian matrix evaluated on a deep learning model to achieve optimal class separability in binary classification tasks. Our approach is substantiated by formal proofs that establish its capability to maximize between-class mean distance (the concept of separation) and minimize within-class variances (the concept of compactness), which together define the two linear discriminant analysis (LDA) criteria, particularly under ideal data conditions such as isotropy around class means and dominant leading eigenvalues. By projecting data into the combined space of the most relevant eigendirections from both matrices, we achieve optimal class separability as per these LDA criteria. Empirical validation across neural and health datasets consistently supports our theoretical framework and demonstrates that our method outperforms established methods. Our method stands out by addressing both separation and compactness criteria, unlike PCA and the Hessian method, which predominantly emphasize one criterion each. This comprehensive approach captures intricate patterns and relationships, enhancing classification performance. Furthermore, through the utilization of both LDA criteria, our method outperforms LDA itself by leveraging higher-dimensional feature spaces, in accordance with Cover’s theorem, which favors linear separability in higher dimensions. Additionally, our approach sheds light on complex DNN decision-making, rendering them comprehensible within a 2D space.
@@ 2.1 Approach: combining eigenanalysis of covariance and Hessian matrices
    Covariance matrix eigenanalysis: A covariance matrix $S(\boldsymbol{\theta})$ is a $D\times D$ matrix, where $D$ represents the dimensionality of the predictor attributes. Each element of the covariance matrix reflects the covariance between the two predictor attributes $\theta_{1}$ and $\theta_{2}$:
    $S(\theta_{1},\theta_{2})=\frac{1}{n-1}\sum_{i=1}^{n}(x_{1i}-\bar{x}_{1})(x_{2i}-\bar{x}_{2})$ (1)
    where $x_{1i}$ and $x_{2i}$ are the corresponding observations of these predictor attributes for the $i$-th data instance, $\bar{x}_{1}$ and $\bar{x}_{2}$ are the sample means of the predictor attributes $\theta_{1}$ and $\theta_{2}$, respectively, and $n$ is the number of data instances or observations.
    Performing eigenanalysis on the covariance matrix $S(\boldsymbol{\theta})$, we obtain eigenvalues $\lambda_{i}$ and corresponding eigenvectors $\mathbf{v}_{i}$. The leading eigenvector $\mathbf{v}_{1}$ associated with the largest eigenvalue $\lambda_{1}$, as captured in the eigen-equation $S(\boldsymbol{\theta})\cdot\mathbf{v}_{1}=\lambda_{1}\cdot\mathbf{v}_{1}$, represents the principal direction with the highest variance.
@@
Theorem 1 (Maximizing covariance for maximizing squared between-class mean distance).
Consider two sets of 1D data points representing two classes, denoted as $C_{1}$ and $C_{2}$, each consisting of $n$ samples. The data in $C_{1}$ and $C_{2}$ are centered around their respective means, $\mu_{1}$ and $\mu_{2}$. Here, $\mu$ denotes the overall mean of the combined data from $C_{1}$ and $C_{2}$. Furthermore, the variances for $C_{1}$ and $C_{2}$ are denoted as $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$, respectively, and the combined data has an overall variance of $\sigma^{2}$. The between-class mean distance, denoted as $d$, represents the separation between the means of $C_{1}$ and $C_{2}$. We establish the following relationship:
$\sigma^{2}=\frac{d^{2}}{4(1-\lambda)},$ (6)
where $\lambda=\frac{1}{2}\left(\frac{\sigma_{1}^{2}}{\sigma^{2}}+\frac{\sigma_{2}^{2}}{\sigma^{2}}\right)$ means a constant between 0 and 1 that reflects the distribution of the original data when $C_{1}$ and $C_{2}$ are considered as projected representations, which implies that maximizing the variance ($\sigma^{2}$) in this context will also maximize the squared between-class mean distance ($d^{2}$).
Proof sketch
@@
    Simplify the expression for the combined data variance $\sigma^{2}$ to $\frac{1}{2}\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right)+\frac{1}{4}d^{2}$.
@@
    Apply the variance ratio preservation property for projection onto a vector to relate $\sigma_{1}^{2}$, $\sigma_{2}^{2}$, and $\sigma^{2}$ through $\lambda$.
@@
In the context of multidimensional data, extending Theorem 1 involves considering data points as $D$-dimensional vectors, where the dataset consists of two subsets, $C_{1}=\{\vec{x}_{i}\in\mathbf{R}^{D}\mid i=1,2,\ldots,n\}$ with covariance $S_{1}$, and $C_{2}=\{\vec{y}_{j}\in\mathbf{R}^{D}\mid j=1,2,\ldots,n\}$ with covariance $S_{2}$. The theorem, however, is supposed to hold for the projection of the data onto an arbitrary 1D subspace of the full D-dimensional data space, and to imply that maximizing $\sigma^{2}$ by choosing the direction of this subspace appropriately will also be the direction that maximizes the difference between the projected sample means. A general case in which it holds is when the individual distributions are isotropic about their means, meaning the covariances $S_{1}$ and $S_{2}$ are proportional to the identity matrix in $\mathrm{D}$ dimensions as mathematically shown in B.
Theorem 2 (Maximizing Hessian for minimizing within-class variance).
@@
$\mathrm{H}_{\theta}=\frac{1}{\sigma_{\text{post}}^{2}},$ (7)
which implies that maximizing the Hessian ($\mathrm{H}_{\theta}$) will minimize the within-class variances ($\sigma_{\text{post}}^{2}$).
Proof sketch
@@ 3.1 General guidelines
      Normalized data: Normalizing the data, typically through techniques such as z-score normalization, ensures that each attribute has a mean of zero and a standard deviation of one. This process equalizes the variance in all directions across both classes, which, as discussed in Subsection 3.2 and formalized in Theorem 3, helps approximate the isotropic condition within each class.
@@
      Low correlation between attributes: Datasets with low correlation between attributes produce covariance matrices with minimized off-diagonal elements. This condition approximates isotropy, which is characterized by covariance matrices being proportional to the identity matrix. Low correlation is also a desirable property of a well-structured dataset, as high correlation (multicollinearity) can undermine the statistical significance of individual variables and reduce the interpretability and robustness of the model [53, 54, 55].
@@
    The presence of a dominant first eigenvalue can be assessed from the eigenspectra of the two matrices, which display the eigenvalues on a log scale. A dominant first eigenvalue is characterized by the eigenvalues being spread over many orders of magnitude, often with approximately uniform spacing between them. When a dominant first eigenvalue is present in the covariance matrix, it suggests a dominant underlying structure or trend in the data. When this occurs in the Hessian matrix, it characterizes a ”sloppy” model [56, 57, 58, 59, 60, 61, 62], which has a few stiff directions (significant eigenvalues) and many sloppy directions (small eigenvalues).
3.2 Z-score normalization for supporting class-level isotropy
Z-score normalization is a standard preprocessing method [63, 64] that is particularly useful when the features in the dataset have different scales or units. By applying z-score normalization, the features are brought onto a common scale, which can improve the performance and stability of machine learning models, especially those that rely on distance calculations or gradient-based optimization [65, 66]. For the proposed method, z-score normalization plays an additional role in promoting isotropic conditions at the class level. While z-score normalization can be applied to each class independently, this approach results in separate feature spaces for each class, making it incompatible for classifying new instances.
The goal of achieving isotropy is to have within-class covariance matrices that approximate a form proportional to the identity matrix. An isotropic covariance matrix is defined by having uniform diagonal elements and minimized off-diagonal elements. While the off-diagonal elements are controlled by maintaining low correlations between attributes, ensuring uniformity (or near uniformity) along the diagonal elements is also crucial for achieving the desired isotropic condition.
The diagonal elements of a within-class covariance matrix correspond to the variances of each individual attribute within the respective class. If the attributes in the original dataset have significantly different scales or units, these variances can differ vastly. In such cases, the covariance matrix would have non-uniform diagonal values, i.e., $\sigma_{w}^{(i)^{2}}\gg\sigma_{w}^{(j)^{2}}$ for some attribute combinations $i,j$ and $w=1,2$, making it far from isotropic. Thus, even with low correlations between attributes, the resulting covariance matrix would not fit the form of a matrix proportional to the identity matrix.
Z-score normalization effectively addresses this issue by transforming each attribute to have a mean of zero and a standard deviation of one across the entire dataset. The following theorem formalizes the impact of z-score normalization on within-class variances:
Theorem 3 (Within-class variance scaling through z-score normalization).
Consider the same setup as in Theorem 1, with two sets of 1D data points representing two classes, denoted as $C_{1}$ and $C_{2}$, each containing $n$ samples. After applying z-score normalization to the entire dataset, the resulting within-class variances, denoted as $\sigma_{1}^{\prime 2}$ and $\sigma_{2}^{\prime 2}$, are given by the following relationships:
$\sigma_{1}^{\prime 2}=\frac{\sigma_{1}^{2}}{\sigma^{2}},\quad\sigma_{2}^{\prime 2}=\frac{\sigma_{2}^{2}}{\sigma^{2}},$
where $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$ are the original within-class variances, and $\sigma^{2}$ is the original overall variance of the combined dataset.
Proof.
Let $\{x_{1},x_{2},\ldots,x_{2n}\}$ represent the original set of 1D data points. As defined in Theorem 1, $C_{1}$ and $C_{2}$ are centered around their respective means, $\mu_{1}$ and $\mu_{2}$, with an overall mean $\mu$ for the combined dataset. The z-scored value for each data point $x_{i}$ is computed as:
$z_{i}=\frac{x_{i}-\mu}{\sigma}.$
Thus, the z-scored dataset is $\{z_{1},z_{2},\ldots,z_{2n}\}$, where each $z_{i}$ represents the standardized value of $x_{i}$.
The within-class variance for the z-scored data for class $C_{1}$ is calculated using the definition of variance:
$\sigma_{1}^{\prime 2}=\frac{1}{n-1}\sum_{i=1}^{n}\left(z_{i}-\bar{z}_{1}\right)^{2},$
where $\bar{z}_{1}$ is the mean of the z-scored data for class $C_{1}$, given by:
$\bar{z}_{1}=\frac{\mu_{1}-\mu}{\sigma}.$
Substituting $z_{i}=\frac{x_{i}-\mu}{\sigma}$, we get:
$\sigma_{1}^{\prime 2}=\frac{1}{n-1}\sum_{i=1}^{n}\left(\frac{x_{i}-\mu}{\sigma}-\frac{\mu_{1}-\mu}{\sigma}\right)^{2}.$
Simplifying the expression, we obtain:
$\sigma_{1}^{\prime 2}=\frac{1}{\sigma^{2}}\cdot\frac{1}{n-1}\sum_{i=1}^{n}\left((x_{i}-\mu_{1})^{2}\right),$
Thus, we have:
$\sigma_{1}^{\prime 2}=\frac{\sigma_{1}^{2}}{\sigma^{2}}.$
Similarly, for class $C_{2}$, the within-class variance for the z-scored data is given by:
$\sigma_{2}^{\prime 2}=\frac{1}{\sigma^{2}}\cdot\frac{1}{n-1}\sum_{j=n+1}^{2n}\left((x_{j}-\mu_{2})^{2}\right)=\frac{\sigma_{2}^{2}}{\sigma^{2}}.$
This completes the proof. ∎
Theorem 3 establishes that after z-score normalization, the within-class variance for a particular attribute is normalized relative to the overall variance of that attribute in the combined dataset. This normalization ensures that the variances for different attributes within the same class become relatively uniform, i.e., $\sigma_{w}^{\prime(i)^{2}}\approx\sigma_{w}^{\prime(j)^{2}}$ for all attribute combinations $i,j$ and $w=1,2$, bounded within $[0,1]$, thereby approximating isotropy within each class.
The degree to which a within-class covariance matrix approximate isotropy depends on how similar the relative variances are across different attributes within the respective class. While perfect uniformity may not be achieved, the application of z-score normalization significantly reduces the disparity between the within-class variances, leading to covariance matrices that better fit the desired isotropic form.
Thus, combining z-score normalization and maintaining low correlations between attributes helps produce within-class covariance matrices that are closer to the ideal isotropic condition, supporting the validity of Theorem 1 in higher dimensions.
@@ 4 Complexity analysis
The overall time complexity of our method encompasses several tasks, including covariance and Hessian matrix computations ($\mathcal{O}(N\cdot F^{2})$), eigenanalysis of these matrices ($\mathcal{O}(F^{3})$), selection of eigenvectors corresponding to leading eigenvalues ($\mathcal{O}(1)$), and data projection into the combined space ($\mathcal{O}(N\cdot F)$). Here, $N$ signifies the number of data points in the training set, while $F$ represents the number of features per data point. It is essential to note that this analysis considers our method’s computational demands under the assumption that a pre-existing DNN is in place.
@@ 5.1 Assessed dimensionality reduction techniques
In our evaluation, we compare nine distinct dimensionality reduction and data projection techniques, each offering unique insights: principal component analysis (PCA), kernel PCA (KPCA) [67], Hessian, uniform manifold approximation and projection (UMAP) [68], locally linear embedding (LLE) [69], linear optimal low-rank projection (LOL) [70], linear discriminant analysis (LDA), kernel discriminant analysis (KDA) [71], and the proposed approach. PCA involves projection onto the primary two covariance eigenvectors, i.e., it applies KPCA with a linear kernel. Similarly, LDA employs KDA with a linear kernel. The Hessian method projects data onto the leading two Hessian eigenvectors, capturing the most significant directions of curvature in the loss landscape. While PCA and the Hessian method rely exclusively on the top two eigenvectors of their respective matrices, the proposed method synergistically integrates the top eigenvector from both the covariance and Hessian matrices, as specified by Eqs(4, 5). Notably, for both KPCA and KDA, the best kernels (linear, polynomial, RBF, sigmoid, or cosine similarity) and their associated parameters (kernel coefficient or degree) are determined using grid search tailored to each dataset.
@@ 6 Illustrative case study: WBCD dataset
Figure 3: Pairplot of the first four predictor attributes in the WBCD dataset, colored by class labels. This pairplot presents only four of the 30 predictor attributes, projecting the 30-dimensional data into 1D and 2D spaces between individual attributes and attribute pairs. The diagonal plots display the distribution of each attribute individually using kernel density estimation (KDE) curves, where separation is assessed based on the overlap of the KDE curves for each class. The non-diagonal plots show scatter plots of pairs of attributes, where separation between classes is evaluated based on the extent of overlap or spread of points corresponding to different class labels. The visualizations show that these projections do not provide good separability between the two classes.
@@
Our proposed method leverages the strengths of both the covariance matrix and the Hessian matrix to achieve optimal class separability. The covariance matrix, computed unsupervisedly from the training set using Eq. 1, is a square and symmetric matrix that captures the statistical spread and relationships between predictor attributes. The Hessian matrix, evaluated on a deep learning model trained on the same training set using Eq. 3, is another square and symmetric matrix that reflects the sensitivity and curvature of the loss function with respect to the model parameters. The DNN test scores are as follows: accuracy: 0.9883, F1 score: 0.9841, AUC ROC: 0.9917, Cohen Kappa score: 0.9749, geometric mean score: 0.9874.
By performing eigenanalysis on the covariance and Hessian matrices, we obtain leading eigenvectors that represent the principal directions of variance and the sharpest curvature, respectively. These eigendirections are crucial for projecting the high-dimensional data into a 2D space that enhances class separability by maximizing the between-class mean distance and minimizing the within-class variances, as described in Theorem 1 and Theorem 2. We project the data into the combined space of the two eigendirections using Eq. 4 and Eq. 5. Figure 4 visually demonstrates the effectiveness of the proposed method by showing the 2D projections of the WBCD dataset. The training set projection highlights the observed visual separation achieved on the training data, while the test set projection confirms good visual separability for new unseen data, underscoring the method’s potential to generalize well. The visual separability observed here will be quantified later using SVM (Figures 6 and 7).
@@
While Figures 4 illustrate the visual separations achieved by the proposed method, Figures 6 and 7 quantify their degrees of separation. These figures present the classification performance using various dimensionality reduction methods, including PCA, the Hessian method, LDA, UMAP, and the proposed method, evaluated with linear SVM on the WBCD dataset. Each subplot shows the separating SVM hyperplane along with the corresponding F-score, providing a quantitative measure of the classification performance. Figure 6 highlights the separation achieved on the training data, where UMAP achieves the highest performance with perfect separation, while the proposed method demonstrates strong performance, surpassing other methods. Figure 7 shows the performance on new unseen test data, where the proposed method outperforms all other methods, including UMAP, demonstrating its superior generalization capability.